# TSD-SR: One-Step Diffusion with Target Score Distillation for Real-World Image Super-Resolution

### v3

@@ 5 Conclusion and Limitation
We propose TSD-SR, an effective and one-step model based on diffusion prior for Real-ISR. TSD-SR utilizes the TSD to enhance the realism of images generated by the distillation model. And it leverages DASM to sample distribution-based samples and accumulate their gradients to enhance the recovery of details. Our experiments have demonstrated that TSD-SR outperforms existing one-step Real-ISR models in both performance and inference speed.
limitations. Although our model boasts excellent inference speed and restoration performance, it still has a large number of model parameters compared to the past GAN/non-diffusion approaches. In the future, we plan to employ pruning or quantization methods to compress the model parameters, striving for a lightweight and efficient Real-ISR model.
@@ Appendix A Comparison with GAN-based Methods
We compared our method with GAN-based approaches in Tab. 7. While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms GAN-based methods on all no-reference metrics. Some researchers have found the limitations of PSNR and SSIM in the field of image super-resolution [61, 55]. The effectiveness of PSNR and SSIM in assessing image fidelity in complex degradation scenarios remains debatable, as pixel-level misalignment often arises when restoring severely degraded images. However, no-reference metrics assess image quality based on the individual image, without the need to forcibly align with the ground truth. Therefore, in more complex and realistic degradation scenarios, no-reference metrics may be more suitable for evaluating the results of image super-resolution. In Appendix C, we further provide the visual comparison between full-reference metrics and human preferences, and in Fig. 9, we present a visual comparison with GAN-based methods. From the visualization, it can be observed that our model achieves better results in terms of texture details compared to GAN methods.
@@ Appendix B More Visual Comparisons
In Figs. 10, 11 and 12, we provide more visual comparisons with other diffusion-based methods. Numerous examples demonstrate the robust restoration capabilities of TSD-SR and the high quality of restored images.
@@ Appendix C Comparisons of Full-reference Metrics and Human Preference
We present additional comparative experiments in Figure 13 to demonstrate that PSNR and SSIM may have limitations in assessing image fidelity under complex degradation scenarios. It can be observed that GAN-based methods with higher PSNR and SSIM produce over-smooth or broken textures, raising concerns about their realism and fidelity. While our approach trades off PSNR and SSIM for natural detail restoration, it achieves enhanced realism and broader perceptual acceptance (Our additional user study reveals that 90.28% of participants prefer ours instead of high PSNR and SSIM methods.).
@@ Appendix D Theory of Target Score Matching
The core idea of Target Score Matching (TSM) is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other. Thus we minimize the MSE loss between the Teacher Model’s predictions of $\mathbf{\hat{z}}_{t}$ and $\mathbf{z}_{t}$ by
$\begin{split}&\mathcal{L}_{\mathrm{MSE}}(\mathbf{\hat{z}},\mathbf{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)\|\epsilon_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-\epsilon_{\psi}(\mathbf{z}_{t};t,c_{y})\|_{2}^{2}\right]\end{split}$ (10)
@@
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{MSE}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)\cdot\underbrace{\frac{\partial{\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})}{\partial\mathbf{\hat{z}}_{t}}}_{\text{Diffusion Jacobian}}\\ &\underbrace{({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y}))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\mathbf{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (11)
where we absorb $\frac{\partial\mathbf{\mathbf{\hat{z}}}_{t}}{\partial\mathbf{\mathbf{\hat{z}}}}$ and the other constant into $w(t)$. The computation of the Diffusion Jacobian term is computationally demanding, as it necessitates backpropagation through the Teacher Model. DreamFusion [36] found that this term struggles with small noise levels due to its training to approximate the scaled Hessian of marginal density. This work also demonstrated that omitting the Diffusion Jacobian term leads to an effective gradient for optimizing. Similar to their approach, we update Eq. 11 by omitting Diffusion Jacobian:
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\\ &\mathbb{E}_{t,\epsilon}\Big{[}w(t)\underbrace{({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};c_{y},t)-{\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\mathbf{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (12)
The effectiveness of the method can be proven by starting from the KL divergence. We can use a Sticking-the-Landing [39] style gradient by thinking of ${\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$. For detailed proof, refer to Appendix 4 of DreamFusion [36]. It demonstrates that the gradient of this loss yields the same updates as optimizing the training loss $\mathcal{L}_{\mathrm{MSE}}$ Eq. 10, excluding the Diffusion Jacobian term.
Compared with the VSD loss, we find that the term “Prediction Residual” has changed, and the two losses are similar in the gradient update mode. Specifically, we find that VSD employs identical inputs for both the Teacher and LoRA models to compute the gradient, while here TSM uses high-quality and suboptimal inputs for the Teacher Model. The losses are related to each other through $\epsilon_{\phi}(\mathbf{\hat{z}}_{t};t,c_{y})$.

### v4

@@ 5 Conclusion and Limitation
We propose TSD-SR, an effective one-step model for Real-ISR based on diffusion priors. TSD-SR utilizes TSD to enhance the realism of images generated by the distillation model, and leverages DASM to sample distribution-based noisy latents and accumulate their gradients, thereby improving detail recovery. Our experiments demonstrate that TSD-SR outperforms existing one-step Real-ISR models in both restoration quality and inference speed.
limitations. Although our model achieves excellent inference speed and restoration performance, it still contains significantly more parameters compared to previous GAN- or non-diffusion-based methods. In future work, we plan to apply pruning or quantization techniques to compress the model, aiming to develop a lightweight and efficient Real-ISR solution.
@@ Appendix A Comparison with GAN-based Methods
We compare our method with GAN-based approaches in Tab. 7. While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms them across all no-reference metrics. Prior studies have highlighted the limitations of PSNR and SSIM for evaluating image super-resolution performance [61, 55]. Their effectiveness in assessing image fidelity in complex degradation scenarios remains debatable, as pixel-level misalignment often arises when restoring severely degraded images. However, no-reference metrics evaluate image quality based solely on the individual image, without requiring alignment with the ground truth. Therefore, in more complex and realistic degradation scenarios, they may offer a more appropriate evaluation of super-resolution results. In Appendix C, we further provide a visual comparison between full-reference metrics and human preferences, and in Fig. 9, we present a visual comparison with GAN-based methods. From these visualizations, it is evident that our model produces more realistic texture details than the GAN-based approaches.
@@ Appendix B More Visual Comparisons
In Figs. 10, 11 and 12, we provide additional visual comparisons with other diffusion-based methods. These examples further demonstrate the robust restoration capabilities of TSD-SR and the high quality of the restored images.
@@ Appendix C Comparisons of Full-reference Metrics and Human Preference
We present additional comparative experiments in Figure 13 to demonstrate the limitations of PSNR and SSIM in assessing image fidelity under complex degradation scenarios. As observed, GAN-based methods with higher PSNR and SSIM scores tend to produce over-smoothed or fragmented textures, raising concerns about their realism and perceptual fidelity. In contrast, our approach sacrifices some PSNR and SSIM performance to achieve more natural detail restoration, resulting in enhanced realism and broader perceptual acceptance. An additional user study shows that 90.28% of participants prefer our results over those of methods with higher PSNR and SSIM scores.
@@ Appendix D Theory of Target Score Matching
The core idea of Target Score Matching (TSM) is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other. Thus, we minimize the MSE loss between the Teacher Model’s predictions on $\boldsymbol{\hat{z}}_{t}$ and $\boldsymbol{z}_{t}$ by
$\begin{split}&\mathcal{L}_{\mathrm{MSE}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)\|\epsilon_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})-\epsilon_{\psi}(\boldsymbol{z}_{t};t,c_{y})\|_{2}^{2}\right]\end{split}$ (10)
@@
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{MSE}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)\cdot\underbrace{\frac{\partial{\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})}{\partial\boldsymbol{\hat{z}}_{t}}}_{\text{Diffusion Jacobian}}\\ &\underbrace{({\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\boldsymbol{z}_{t};t,c_{y}))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\boldsymbol{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (11)
where we absorb $\frac{\partial\boldsymbol{\boldsymbol{\hat{z}}}_{t}}{\partial\boldsymbol{\boldsymbol{\hat{z}}}}$ and the other constant into $w(t)$. The computation of the Diffusion Jacobian term is computationally demanding, as it necessitates backpropagation through the Teacher Model. DreamFusion [36] found that this term struggles with small noise levels due to its training to approximate the scaled Hessian of marginal density. This work also demonstrated that omitting the Diffusion Jacobian term leads to an effective gradient for optimizing. Similar to their approach, we update Eq. 11 by omitting Diffusion Jacobian:
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})=\\ &\mathbb{E}_{t,\epsilon}\Big{[}w(t)\underbrace{({\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};c_{y},t)-{\epsilon}_{\psi}(\boldsymbol{z}_{t};c_{y},t))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\boldsymbol{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (12)
The effectiveness of the method can be proven by starting from the KL divergence. We can use a Sticking-the-Landing [39] style gradient by thinking of ${\epsilon}_{\psi}(\boldsymbol{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$. For detailed proof, refer to Appendix 4 of DreamFusion [36]. It demonstrates that the gradient of this loss yields the same updates as optimizing the training loss $\mathcal{L}_{\mathrm{MSE}}$ Eq. 10, excluding the Diffusion Jacobian term.
Compared with the VSD loss, we find that the term “Prediction Residual” has changed, and the two losses are similar in the gradient update mode. Specifically, we find that VSD employs identical inputs for both the Teacher and LoRA models to compute the gradient, while here TSM uses high-quality and suboptimal inputs for the Teacher Model. The losses are related to each other through $\epsilon_{\phi}(\boldsymbol{\hat{z}}_{t};t,c_{y})$.